OpenStudy (anonymous):

The sum of three consecutive odd integers is 57. What are the three numbers? A. 17, 19, 21 B. 15, 19, 23 C. 19, 21, 23 D. 18, 19, 20

OpenStudy (austinl):

They are all integers, however, they have to be odd, so that removes D. They have to be consecutive, so that removes B. To find your answer, just add up the numbers in A and C and see what you get for answers!

OpenStudy (texaschic101):

x = first number x + 2 = second number x + 4 = third number x + x + 2 + x + 4 = 57 combine like terms and solve for x.(your smallest number) Once x is found, sub that back into the x + 2 and x + 4 to find those values.

OpenStudy (austinl):

That is a great method, I must admit. Probably the better answer because it should work if you aren't given answer choices.

OpenStudy (texaschic101):

I like your elimination method :)

OpenStudy (tkhunny):

Keep your mind open to other ways to make your life easier. If you define: x, x+2 and x+4, solving for x produces the lowest value. That's fine. If you define: x - 2, x and x + 2, solving for x produces the middle value. That's fine and the addition might be easier. However, it is important to note that neither of those definitions will necessarily produce an ODD result. In this case, the sum is 51 and we have only addition, so it turns out okay. You might want to give some thought how you could enforce ODD and give it another go.

OpenStudy (anonymous):

Thanks guys i got A.